OpenStudy (anonymous):

Which of the following is the equation of a line perpendicular to the line y=-1/3 x+1, passing through the point (2,7)? A. -3x - y = 1 B. -3x + y = -1 C. 3x - y = 1 D. 3x - y = -1

OpenStudy (anonymous):

slope of perpendicular line is m=3\[y-7=3(x-2)\]\[y=3x-6+7\]\[-3x+y=1\]\[3x-y=-1\]answer D